### v3

@@ Appendix B The other TNG50 and Auriga galaxies located within $3\sigma$ confidence level
Despite the four galaxies with DM halos and satellite planes vertically aligned due to disk tilt, there are another nine galaxies, TNG50 392277, 402555, 530852, 539333, 550149, 550475, 566365, 571454, and 580406 with DM shape outside of the $1\sigma$ but within the $3\sigma$ confidence level of the Milky Way. These galaxies have intrinsically near-spherical DM halos. Four of them, TNG50 530852, 550149, 571454, 580406 also experienced significant disk tilt with tilt angles from $30^{\circ}$ to $100^{\circ}$. The disk tilts of the former three galaxies are associated with and likely induced by a massive minor merger. Such events occurred quite often and were not necessarily associated with systems exhibiting satellite planes. The disk of TNG50 580406 was tilted about $30^{\circ}$ and was not associated with any significant mergers.

### v4

@@ Appendix B The other TNG50 and Auriga galaxies located within $3\sigma$ confidence level
Despite the four galaxies with DM halos and satellite planes vertically aligned due to disc tilt, there are another nine galaxies, TNG50 392277, 402555, 530852, 539333, 550149, 550475, 566365, 571454, and 580406 with DM shape outside of the $1\sigma$ but within the $3\sigma$ confidence level of the Milky Way. These galaxies have intrinsically near-spherical DM halos. Four of them, TNG50 530852, 550149, 571454, 580406 also experienced significant disc tilt with tilt angles from $30^{\circ}$ to $100^{\circ}$. The disc tilts of the former three galaxies are associated with and likely induced by a massive minor merger. Such events occurred quite often and were not necessarily associated with systems exhibiting satellite planes. The disc of TNG50 580406 was tilted about $30^{\circ}$ and was not associated with any significant mergers.